# Towards a theory of dissipative Dark Matter I: the Born limit

### v1

@@ D.2 Vector emission from fermionic DM candidates
by assuming that, for a massless emitted particle, the polarization vectors are in the Coulomb gauge, $\epsilon(q)=(0,\mathbf{e})$, where $\mathbf{e}$ denotes the transverse polarization vector in Cartesian components satisfying $\mathbf{e}\cdot\mathbf{q}=0$. The spinor expansion separates into $C^{(\prime)}$ terms, which are independent of Pauli matrices, and $D^{(\prime)}$ terms, which explicitly depend on them. However, both for dipole and quadrupole emission, only the $C^{(\prime)}$-terms are relevant compared to the $D^{(\prime)}$-term. Consequently, (D.12) is spin-independent for both types of emission.
@@
$=\frac{-4g_{1}^{2}g_{2}m_{2}}{\omega|\vec{p}_{i}-\vec{p}_{f}|^{2}}\vec{q}\cdot\vec{e}\,\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}}=0,$
@@
$(\vec{e}\cdot\bm{\sigma})(\vec{q}\cdot\bm{\sigma})=i(\vec{e}\times\vec{q})\cdot\bm{\sigma},\quad(\vec{q}\cdot\bm{\sigma})(\vec{e}\cdot\bm{\sigma})=i(\vec{q}\times\vec{e})\cdot\bm{\sigma},$ (D.15)
which, in the classical theory of electromagnetism and QED, give the direction of the magnetic field. The cancellation between these two terms at quadrupole order, as explicitly shown in (D.14), is consistent with the vanishing classical magnetic dipole of a two-particle system [96]. Accordingly, the processes $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}V$ and $S_{1}S_{2}\to S_{1}S_{2}V$ yield the same dipole and quadrupole emission amplitudes in the case of distinguishable particle self-scattering. For identical particle self-scattering, both the $t$- and $u$-channels must be included according to the underlying statistical properties of the DM candidates, leading to different emission amplitudes in the two dissipative scenarios.
Beyond the quadrupole order, the denominator expansion yields different kinematic prefactors for the two components defined in (D.15), implying that spin-dependent effects only arise at higher orders. The subleading nature of these effects is expected in a QED-like model, where magnetic transitions are known to be suppressed compared to electric ones [61].
@@ Appendix E Derivation of $\mathbf{q}$-dependent eikonal emission factors
On the other hand, the soft factors (4.1) and (4.3) scale as
@@ E.1 Vector soft emission
$\mathcal{E}_{V}=g_{1}\frac{\epsilon^{*}(q)\cdot p_{3}}{p_{3}\cdot q}+g_{2}\frac{\epsilon^{*}(q)\cdot p_{4}}{p_{4}\cdot q}-g_{2}\frac{\epsilon^{*}(q)\cdot p_{2}}{p_{2}\cdot q}-g_{1}\frac{\epsilon^{*}(q)\cdot p_{1}}{p_{1}\cdot q}.$ (E.3)
We make the assumption of a massless emitted vector boson in Coulomb gauge $\epsilon(q)=(0,\mathbf{e})$, where $\mathbf{e}$ denotes the transverse polarization vector in Cartesian components, and obtain for the $v_{i}$-leading order of (E.3)
$\mathcal{E}_{V}$ $\overset{CM}{\simeq}\frac{1}{\omega}\left[\frac{g_{1}}{m_{1}}(\vec{p}_{f}-\vec{p}_{i})-\frac{g_{2}}{m_{2}}(\vec{p}_{f}-\vec{p}_{i})\right]\cdot\textbf{e}+\mathcal{O}(v_{i}^{0})$ (E.4)
$\simeq\frac{1}{\omega}\left(\frac{g_{1}}{m_{1}}-\frac{g_{2}}{m_{2}}\right)(\vec{p}_{f}-\vec{p}_{i})\cdot\textbf{e}+\mathcal{O}(v_{i}^{0}),$
@@
[…]
$\mathcal{E}_{V}|_{\chi_{2}\text{ outgoing}}$ $=g_{2}\frac{p_{4}\cdot\epsilon(q)}{p_{4}\cdot q}\frac{k_{\text{elastic}}^{2}}{k_{4}^{2}}\overset{\text{CM}}{\simeq}g_{2}\frac{-\mathbf{p}_{f}\cdot\mathbf{e}}{m_{2}\omega+\mathbf{p}_{f}\cdot\mathbf{q}+\mathcal{O}(v_{i}^{4})}\frac{|\vec{p}_{i}-\vec{p}_{f}|^{2}}{|\vec{p}_{i}-\vec{p}_{f}-{\vec{q}\mu}/{m_{2}}|^{2}}.$ (E.6)
During the $v_{i}$-expansion of (E.6), the contribution from the correction factor (4.21) is trivial at leading order and only appears at subleading order. It introduces some $\mathbf{q}$-dependent terms necessary to ensure consistency between the soft energy differential cross-section and the full energy differential cross-section over a significant energy range.
@@ E.2 Scalar soft emission
$\mathcal{E}_{\phi\chi}=\frac{y_{1}m_{1}}{p_{3}\cdot q}+\frac{y_{2}m_{2}}{p_{4}\cdot q}-\frac{y_{2}m_{2}}{p_{2}\cdot q}-\frac{y_{1}m_{1}}{p_{1}\cdot q}-\frac{2y_{1}}{m_{1}}-\frac{2y_{2}}{m_{2}}.$ (E.7)
@@ F.1 Massive vector boson emission
$\varepsilon^{\mu}_{(L)}=\frac{1}{m_{V}}\left(|\mathbf{q}|,\,\omega\frac{\mathbf{q}}{|\mathbf{q}|}\right).$ (F.1)
@@
$\varepsilon^{\mu}_{(L)}\underset{m_{V}\rightarrow 0}{=}\frac{q^{\mu}}{m_{V}}+\mathcal{O}\left(\frac{m_{V}}{\omega}\right),$ (F.2)
@@
$\mathcal{M}^{(L)}=\varepsilon^{(L)}_{\mu}\mathcal{M}^{\mu}=-\frac{m_{V}}{\omega}\hat{{q}}^{i}\mathcal{M}^{i}=-\frac{m_{V}}{|\mathbf{q}|}\mathcal{M}^{0}.$ (F.4)
@@
F.2 Emission amplitudes and energy differential cross-sections
In the following, we present the corresponding $m_{\phi,V}$-corrections to the emission amplitudes and the energy differential cross-sections for distinguishable particle scattering.
The squared amplitudes are obtained, for the transverse modes of $V$-emission, by first summing over the transverse polarization states $\lambda$ of the polarization vector $\mathbf{e}^{(\lambda)}$ and then averaging over the direction of the emitted particle, whereas for $\phi$-emission and the longitudinal $V$-emission only the angular averaging is required. Accordingly, we use the identities given in (3.28).
@@ F.2.1 Long-range mediation
$\mathcal{M}_{\mathcal{D}}^{(L)}=\varepsilon_{\nu}^{(L)}\mathcal{M_{D}}^{\nu}=4\mathcal{F_{D}}\frac{m_{V}}{\omega}\frac{m_{1}m_{2}(\vec{v}_{f}-\vec{v}_{i})\cdot\vec{\bm{\hat{}}{\mathbf{q}}}}{\mu\omega|\vec{v}_{i}-\vec{v}_{f}|^{2}}\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}},$ (F.5)
@@
$\mathcal{M}_{\mathcal{D}}^{(T)}=\varepsilon_{\nu}^{(T)}\mathcal{M_{D}}^{\nu}=4\mathcal{F_{D}}\frac{m_{1}m_{2}(\vec{v}_{f}-\vec{v}_{i})\cdot\vec{e}}{\mu\omega|\vec{v}_{i}-\vec{v}_{f}|^{2}}\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}}.$ (F.6)
We emphasize that, at this order, the spinor expansion yields only trivial contributions. As a result, for the process $S_{1}S_{2}\to S_{1}S_{2}\,V$, the dipole components of the emission amplitude are also given by (F.5) and (F.6), with the spinor structures reducing to $\delta_{s_{1}s_{3}}\,\delta_{s_{2}s_{4}}\to 1$. Upon squaring these transverse and longitudinal components and performing the angular average individually using (3.17), and subsequently summing these contributions, the resulting expression can be inserted into (3.13) to obtain the energy-differential cross section (3.22).
@@
$+\left.\left.\left.\hat{{v}}_{i}^{j}|\mathbf{v}_{f}||\mathbf{v}_{i}|(|\mathbf{v}_{i}|^{2}-|\mathbf{v}_{f}|^{2})\right)-(\mathbf{v}_{i}\leftrightarrow\mathbf{v}_{f})\right)\right.\left.+\frac{\delta^{jk}}{2\mu}|\vec{v}_{i}-\vec{v}_{f}|^{2}\omega\right]\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}},$
@@
We note that the $m_{V}$-dependence of (F.8) appears from the on-shell relation of the emitted particle, i.e., $|\vec{q}|=\sqrt{\omega^{2}-m_{V}^{2}}$. In both components, we have simplified the $\vec{e}\cdot\vec{\hat{q}}$ terms arising from the spinor expansion. Consequently, the components of the quadrupole emission amplitudes for scalar DM candidates are also given by (F.7) and (F.8), treating the spinor factors as trivial.
For the polarization modes given in (F.7) and (F.8) individually, we square the corresponding amplitudes, sum over the final spins, average over the initial ones, and perform the angular average using (3.28). The sum over the three polarization states is then carried out by adding the two resulting contributions, and the combined expression is finally inserted into (3.13) to obtain the energy-differential cross section,
@@
For scalar DM candidates, the dipole emission amplitude coincides with (F.10), up to the replacement $\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}}\to 1$. This emission amplitude, squared and angular-averaged using (3.17), and subsequently inserted into (3.13), yields the energy differential cross-section (3.22), which also applies to fermionic DM candidates.
@@
$\left.\left.\left.+\hat{{v}}_{i}^{j}|\mathbf{v}_{f}||\mathbf{v}_{i}|(|\mathbf{v}_{i}|^{2}-|\mathbf{v}_{f}|^{2})\right)-(\mathbf{v}_{i}\leftrightarrow\mathbf{v}_{f})\right)\right.\left.+\frac{\delta^{jk}}{\mu}|\vec{v}_{i}-\vec{v}_{f}|^{2}\omega\right.\Big{]}\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}},$
@@
$+\left.\left.\left.\hat{{v}}_{i}^{j}|\mathbf{v}_{f}||\mathbf{v}_{i}|(|\mathbf{v}_{i}|^{2}-|\mathbf{v}_{f}|^{2})\right)-(\mathbf{v}_{i}\leftrightarrow\mathbf{v}_{f})\right)\right.+\left.\frac{\delta^{jk}}{2\mu}|\vec{v}_{i}-\vec{v}_{f}|^{2}\omega\right.\Big{]}\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}}.$
@@ F.2.2 Short-range mediated force
$\mathcal{M}^{(T)}_{\mathcal{Q}}=4\mathcal{F^{\prime}_{Q}}\frac{m_{1}m_{2}\mu^{2}|\vec{q}|}{m_{V^{\prime}}^{2}\omega^{2}}e^{j}\hat{q}^{k}\left[{|\mathbf{v}_{f}|^{2}\hat{v}_{f}^{j}\hat{v}_{f}^{k}-|\mathbf{v}_{i}|^{2}\hat{v}_{i}^{j}\hat{v}_{i}^{k}}\right]\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}}.$ (F.18)
Similarly to long-range mediation, the $m_{V}$-dependency of (F.18) only comes from the on-shell mass relation of the emitted particle in $|\mathbf{q}|$. Moreover, we have simplified the $\vec{e}\cdot\vec{\hat{q}}$ terms in (F.17) and (F.18) arising from the spinor expansion. Accordingly, the quadrupole emission amplitude apply equally well to scalar DM candidates, with the spinor Kronecker deltas being replaced by unity.
Each polarization component is squared, summed over the final spins, averaged over the initial ones, and angularly averaged over the emission direction using (3.28). The resulting contributions are then summed to obtain the squared amplitude summed over the three polarization states. Once the corresponding expression is inserted into (3.13), we obtain the energy-differential cross section,
@@
For the scalar emission amplitude associated at sub-leading order in $r$, we follow the same decomposition as in (F.11), analogously to the short-range mediation case. In the case of fermionic DM candidates, the piece of the emission amplitude that arises from massless emission reads
@@
From these matrix elements, by adding them according to (F.11) and squaring them, summing over the initial spins, averaging over the final ones, and averaging over the direction of the emitted particle using (3.28), we obtain the energy differential cross-section
@@
The corresponding energy differential cross-section, obtained from the sum of these matrix elements according to (F.11), squared, and averaged over the direction of the emitted particle using (3.28), reads

### v2

@@ D.2 Vector emission from fermionic DM candidates
by assuming that, for a massless emitted particle, the polarization vectors are in the Coulomb gauge, $\varepsilon^{\mu*}(q)=(0,\mathbf{e}^{*})$, where $\mathbf{e}^{*}$ denotes the transverse polarization vector in Cartesian components satisfying $\mathbf{e}^{*}\cdot\mathbf{q}=0$. The spinor expansion separates into $C^{(\prime)}$ terms, which are independent of Pauli matrices, and $D^{(\prime)}$ terms, which explicitly depend on them. However, both for dipole and quadrupole emission, only the $C^{(\prime)}$-terms are relevant compared to the $D^{(\prime)}$-term. Consequently, (D.12) is spin-independent for both types of emission.
@@
$=\frac{-4g_{1}^{2}g_{2}m_{2}}{\omega|\vec{p}_{i}-\vec{p}_{f}|^{2}}\vec{q}\cdot\vec{e}^{*}\,\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}}=0,$
@@
$(\vec{e}^{*}\cdot\bm{\sigma})(\vec{q}\cdot\bm{\sigma})=i(\vec{e}^{*}\times\vec{q})\cdot\bm{\sigma},\quad(\vec{q}\cdot\bm{\sigma})(\vec{e}^{*}\cdot\bm{\sigma})=i(\vec{q}\times\vec{e}^{*})\cdot\bm{\sigma},$ (D.15)
which, in QED, corresponds to magnetic-like components. Beyond the quadrupole order, the denominator expansion yields different kinematic prefactors for the two components defined in (D.15), implying that spin-dependent effects only arise at higher orders. The subleading nature of these effects is expected in a QED-like model, where spin-induced magnetic transitions are known to be suppressed compared to electric ones [73]. Accordingly, the processes $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}V$ and $S_{1}S_{2}\to S_{1}S_{2}V$ yield the same dipole and quadrupole emission amplitudes in the case of distinguishable particle self-scattering. For identical particle self-scattering, both the $t$- and $u$-channels must be included according to the underlying statistical properties of the DM candidates, leading to different emission amplitudes in the two dissipative scenarios.
@@ Appendix E Derivation of $\mathbf{q}$-dependent eikonal emission factors
On the other hand, the soft factors (4.1), (4.2) and (4.3) scale as
@@ E.1 Vector soft emission
$\mathcal{E}_{V}=g_{1}\frac{\varepsilon^{*}(q)\cdot p_{3}}{p_{3}\cdot q}+g_{2}\frac{\varepsilon^{*}(q)\cdot p_{4}}{p_{4}\cdot q}-g_{2}\frac{\varepsilon^{*}(q)\cdot p_{2}}{p_{2}\cdot q}-g_{1}\frac{\varepsilon^{*}(q)\cdot p_{1}}{p_{1}\cdot q}.$ (E.3)
We make the assumption of a massless emitted vector boson in Coulomb gauge $\varepsilon^{\mu*}(q)=(0,\mathbf{e}^{*})$, where $\mathbf{e}^{*}$ denotes the transverse polarization vector in Cartesian components, and obtain for the $v_{i}$-leading order of (E.3)
$\mathcal{E}_{V}$ $\overset{CM}{\simeq}-\frac{1}{\omega}\left[\frac{g_{1}}{m_{1}}(\vec{p}_{f}-\vec{p}_{i})-\frac{g_{2}}{m_{2}}(\vec{p}_{f}-\vec{p}_{i})\right]\cdot\textbf{e}^{*}+\mathcal{O}(v_{i}^{0})$ (E.4)
$\simeq-\frac{1}{\omega}\left(\frac{g_{1}}{m_{1}}-\frac{g_{2}}{m_{2}}\right)(\vec{p}_{f}-\vec{p}_{i})\cdot\textbf{e}^{*}+\mathcal{O}(v_{i}^{0}),$
@@
In the case of processes with an equal coupling-to-mass ratio, this term vanishes, and $v_{i}$-subleading terms of (E.3) become the dominant terms. The corresponding derivation is similar, but the inclusion of the correction factor (4.21) is necessary to obtain (4.16). The relevance of this factor in the quadrupole soft factor becomes apparent when considering, for example, an emission from the final external $\chi_{2}$-leg,
$\mathcal{E}_{V}|_{\chi_{2}\text{ outgoing}}$ $=g_{2}\frac{p_{4}\cdot\varepsilon^{*}(q)}{p_{4}\cdot q}\frac{k_{\text{elastic}}^{2}}{k_{4}^{2}}\overset{\text{CM}}{\simeq}g_{2}\frac{\mathbf{p}_{f}\cdot\mathbf{e}^{*}}{m_{2}\omega+\mathbf{p}_{f}\cdot\mathbf{q}+\mathcal{O}(v_{i}^{4})}\frac{|\vec{p}_{i}-\vec{p}_{f}|^{2}}{|\vec{p}_{i}-\vec{p}_{f}-{\vec{q}\mu}/{m_{2}}|^{2}}.$ (E.6)
During the $v_{i}$-expansion of (E.6), the contribution from the correction factor (4.21) is trivial at leading order and only appears at subleading order. It introduces some $\mathbf{q}$-dependent terms necessary to ensure consistency between the soft energy-differential cross section and the full energy-differential cross section over a significant energy range.
@@ E.2 Scalar soft emission
$\mathcal{E}_{\phi\chi}=\frac{y_{1}m_{1}}{p_{3}\cdot q}+\frac{y_{2}m_{2}}{p_{4}\cdot q}-\frac{y_{2}m_{2}}{p_{2}\cdot q}-\frac{y_{1}m_{1}}{p_{1}\cdot q}+\frac{y_{1}}{m_{1}}+\frac{y_{2}}{m_{2}}.$ (E.7)
@@ F.1 Massive vector boson emission
$\varepsilon^{\mu*}_{(L)}=\frac{1}{m_{V}}\left(|\mathbf{q}|,\,\omega\frac{\mathbf{q}}{|\mathbf{q}|}\right).$ (F.1)
@@
$\varepsilon^{\mu*}_{(L)}\underset{m_{V}\rightarrow 0}{=}\frac{q_{\mu}}{m_{V}}+\mathcal{O}\left(\frac{m_{V}}{\omega}\right),$ (F.2)
@@
$\mathcal{M}^{(L)}=\varepsilon^{(L)*}_{\mu}\mathcal{M}^{\mu}=-\frac{m_{V}}{\omega}\hat{{q}}^{i}\mathcal{M}^{i}=-\frac{m_{V}}{|\mathbf{q}|}\mathcal{M}^{0}.$ (F.4)
@@
F.2 Emission amplitudes and energy-differential cross sections
In the following, we present the corresponding $m_{\phi,V}$-corrections to the emission amplitudes and the energy-differential cross sections for distinguishable particle scattering.
The squared amplitudes are obtained, for the transverse modes of $V$-emission, by first summing over the transverse polarization states $\lambda$ of the polarization vector $\mathbf{e}^{(\lambda)*}$ and then averaging over the direction of the emitted particle, whereas for $\phi$-emission and the longitudinal $V$-emission only the angular averaging is required. Accordingly, we use the identities given in (3.13), (3.17) and (3.23)–(3.24).
@@ F.2.1 Long-range mediation
$\mathcal{M}_{\mathcal{D}}^{(L)}=\varepsilon_{\nu}^{(L)*}\mathcal{M_{D}}^{\nu}=4\mathcal{F_{D}}\frac{m_{V}}{\omega}\frac{m_{1}m_{2}(\vec{v}_{f}-\vec{v}_{i})\cdot\vec{\bm{\hat{}}{\mathbf{q}}}}{\mu\omega|\vec{v}_{i}-\vec{v}_{f}|^{2}}\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}},$ (F.5)
@@
$\mathcal{M}_{\mathcal{D}}^{(T)}=\varepsilon_{\nu}^{(T)*}\mathcal{M_{D}}^{\nu}=4\mathcal{F_{D}}\frac{m_{1}m_{2}(\vec{v}_{f}-\vec{v}_{i})\cdot\vec{e}^{*}}{\mu\omega|\vec{v}_{i}-\vec{v}_{f}|^{2}}\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}}.$ (F.6)
We emphasize that, at this order, the spinor expansion yields only trivial contributions. As a result, for the process $S_{1}S_{2}\to S_{1}S_{2}\,V$, the dipole components of the emission amplitude are also given by (F.5) and (F.6), with the spinor structures reducing to $\delta_{s_{1}s_{3}}\,\delta_{s_{2}s_{4}}\to 1$. Upon squaring these transverse and longitudinal components and performing the angular average individually using (3.23)–(3.24), and subsequently summing these contributions, the resulting expression can be inserted into (3.19) to obtain the energy-differential cross section (3.28).
@@
$+\left.\left.\left.\frac{1}{2}\hat{{v}}_{i}^{j}|\mathbf{v}_{f}||\mathbf{v}_{i}|(|\mathbf{v}_{i}|^{2}-|\mathbf{v}_{f}|^{2})\right)-(\mathbf{v}_{i}\leftrightarrow\mathbf{v}_{f})\right)\right.\left.+\frac{\delta^{jk}}{2\mu}|\vec{v}_{i}-\vec{v}_{f}|^{2}\omega\right]\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}},$
@@
We note that the $m_{V}$-dependence of (F.8) appears from the on-shell relation of the emitted particle, i.e., $|\vec{q}|=\sqrt{\omega^{2}-m_{V}^{2}}$. Moreover, we have simplified the $\vec{e}^{*}\cdot\vec{\hat{q}}$ terms arising from the spinor expansion. Consequently, the components of the quadrupole emission amplitudes for scalar DM candidates are also given by (F.7) and (F.8), treating the spinor factors as trivial.
For the polarization modes given in (F.7) and (F.8) individually, we square the corresponding amplitudes, sum over the final spins, average over the initial ones, and perform the angular average using (3.13) and (3.17). The sum over the three polarization states is then carried out by adding the two resulting contributions, and the combined expression is finally inserted into (3.19) to obtain the energy-differential cross section,
@@
For scalar DM candidates, the dipole emission amplitude coincides with (F.10), up to the replacement $\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}}\to 1$. This emission amplitude, squared and angular-averaged using (3.23), and subsequently inserted into (3.19), yields the energy-differential cross section (3.28), which also applies to fermionic DM candidates.
@@
$\left.\left.\left.+\frac{1}{2}\hat{{v}}_{i}^{j}|\mathbf{v}_{f}||\mathbf{v}_{i}|(|\mathbf{v}_{i}|^{2}-|\mathbf{v}_{f}|^{2})\right)-(\mathbf{v}_{i}\leftrightarrow\mathbf{v}_{f})\right)\right.\left.+\frac{\delta^{jk}}{\mu}|\vec{v}_{i}-\vec{v}_{f}|^{2}\omega\right.\Big]\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}},$
@@
$+\left.\left.\left.\frac{1}{2}\hat{{v}}_{i}^{j}|\mathbf{v}_{f}||\mathbf{v}_{i}|(|\mathbf{v}_{i}|^{2}-|\mathbf{v}_{f}|^{2})\right)-(\mathbf{v}_{i}\leftrightarrow\mathbf{v}_{f})\right)\right.+\left.\frac{\delta^{jk}}{2\mu}|\vec{v}_{i}-\vec{v}_{f}|^{2}\omega\right.\Big]\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}}.$
@@ F.2.2 Short-range mediated force
$\mathcal{M}^{(T)}_{\mathcal{Q}}=4\mathcal{F^{\prime}_{Q}}\frac{m_{1}m_{2}\mu^{2}|\vec{q}|}{m_{V^{\prime}}^{2}\omega^{2}}e^{*j}\hat{q}^{k}\left[{|\mathbf{v}_{f}|^{2}\hat{v}_{f}^{j}\hat{v}_{f}^{k}-|\mathbf{v}_{i}|^{2}\hat{v}_{i}^{j}\hat{v}_{i}^{k}}\right]\delta_{s_{1}s_{3}}\delta_{s_{2}s_{4}}.$ (F.18)
Similarly to long-range mediation, the $m_{V}$-dependency of (F.18) only comes from the on-shell mass relation of the emitted particle in $|\mathbf{q}|$. Moreover, we have simplified the $\vec{e}^{*}\cdot\vec{\hat{q}}$ terms arising from the spinor expansion. Accordingly, the quadrupole emission amplitudes apply equally well to scalar DM candidates, with the spinor Kronecker deltas being replaced by unity.
Each polarization component is squared, summed over the final spins, averaged over the initial ones, and angularly averaged over the emission direction using (3.13) and (3.17). The resulting contributions are then summed to obtain the squared amplitude summed over the three polarization states. Once the corresponding expression is inserted into (3.19), we obtain the energy-differential cross section,
@@
For the scalar emission amplitude associated at sub-leading order in $r$, we follow the same decomposition as in (F.11), analogously to the long-range mediation case. In the case of fermionic DM candidates, the piece of the emission amplitude that arises from massless emission reads
@@
From these matrix elements, by adding them according to (F.11) and squaring them, summing over the initial spins, averaging over the final ones, and averaging over the direction of the emitted particle using (3.17), we obtain the energy-differential cross section
@@
The corresponding energy-differential cross section, obtained from the sum of these matrix elements according to (F.11), squared, and averaged over the direction of the emitted particle using (3.17), reads